OpenStudy (anonymous):

solve the following by substitution:3x-y=7 5x+2y=8

OpenStudy (anonymous):

3x-y=7 ==> 3x-7=y (1) 5x+2y=8 =(1)=> 5x+2(3x-7)=8 ==> 5x +6x=8-14 ==> 11x=-6 y=3(-11/6) -7